# A Theory of the NEPv Approach for Optimization On the Stiefel Manifold

### v2

@@ 6.1 The NEPv Ansatz
    It is tempting to stipulate $f(\widehat{P})\geq f(P)+\omega\eta$, but that is either false or just hard to prove for the one in Example 6.1 and some of those in Table 1 that involve $D$. Often in our algorithms to solve (1.1) iteratively, with $P$ being the current approximate maximizer, assuming the NEPv Ansatz, we naturally compute $\widehat{P}$ that maximizes $\operatorname{tr}(X^{\operatorname{T}}H(P)X)$ over $X\in\mathbb{O}^{n\times k}$. With that $\widehat{P}$, settling whether $f(\widehat{P})\geq f(P)+\omega\eta$ or not can be a hard task, as in Example 6.1 where the objective function involves $\operatorname{tr}((P^{\operatorname{T}}D)^{2})$.
@@
$H(P)$ conditions $\mathbb{P}$ by SEP $A$ none $\mathbb{O}^{n\times k}$ [20], [29, p.248] OLDA (6.4) with $\theta=1$, $D=0$ $B\succeq 0$, $s_{k}(B)>0$ $\mathbb{O}^{n\times k}$ (6.11) OCCA (6.4) with $\theta=1/2$, $A=0$ $B\succeq 0$, $s_{k}(B)>0$ $\mathbb{O}_{D+}^{n\times k}$ (6.11) $\Theta$TR (6.4) for case $10<\theta<$ 1 $B\succeq 0$, $s_{k}(B)>0$ $\mathbb{O}_{D+}^{n\times k}$ (6.11) $\operatorname{tr}(P^{\operatorname{T}}AP+P^{\operatorname{T}}D)\geq 0$ (6.4) for case $\theta=1$ $B\succeq 0$, $s_{k}(B)>0$ $\mathbb{O}_{D+}^{n\times k}$ (6.11) MBSub (6.4) with $\theta=0$ none $\mathbb{O}_{D+}^{n\times k}$ (6.11) SumCT (6.5) $A_{i}\succeq 0\,\forall i$ [66] Thm. 8.3 TrCP $2\sum_{i=1}^{N}\phi_{i}(\boldsymbol{x})A_{i}$ convex $\phi$ $\mathbb{O}^{n\times k}$ Expl. 8.1 UMDS $4\sum_{i=1}^{N}A_{i}PP^{\operatorname{T}}A_{i}$ $A_{i}\succeq 0\,\forall i$ $\mathbb{O}^{n\times k}$ Expl. 8.2 DFT $2A+2\sum_{i=1}^{n}\phi_{i}(\boldsymbol{x})\boldsymbol{e}_{i}\boldsymbol{e}_{i}^{\operatorname{T}}$ convex $\phi$ $\mathbb{O}^{n\times k}$ Expl. 8.1 * $\phi_{i}(\boldsymbol{x}):=\partial\phi(\boldsymbol{x})/\partial x_{i}$ for $\boldsymbol{x}=[x_{i}]$. $\Theta$TR for $\theta=0$ becomes MBSub.
As to the validity of the NEPv Ansatz on the objective functions in Table 1, it holds for all, except SumTR, under mild conditions on the constant matrices and function $\phi$. Table 4 provides the details on $H(P)$ and conditions under which the NEPv Ansatz holds, where the last column refers to places for justifications. We leave $\mathbb{P}$ unspecified for SumCT but refer it to [66] because it is more complicated to fit the space in the table. In fact, it is required that each $P_{i}$ falls in $\{X\in\mathbb{O}^{n\times k_{i}}\,:\,X^{\operatorname{T}}D_{i}\succeq 0\}$. Later in subsection 8.2 we will argue that for SumCT it would be more efficient to go for the NPDo approach in Part I. For SumTR with $H(P)$ given by (6.2), $\mathbb{P}=\mathbb{O}^{n\times k}$ and $Q=I_{k}$ and the NEPv Ansatz does not hold. This can be drawn from the counterexample, [76, Example 4.1], for which SCF diverges, but later we will show, under the NEPv Ansatz, SCF is guaranteed to converge!
@@
Let function $f$ be defined on some neighborhood $\mathbb{O}_{\delta}^{n\times k}$ of $\mathbb{O}^{n\times k}$ and let $H(P)$ be as in (6.5). Then the NPDo Ansatz implies the NEPv Ansatz.
@@
Suppose that the NPDo Ansatz holds. Given $\widehat{P}\in\mathbb{O}^{n\times k},\,P\in\mathbb{P}\subseteq\mathbb{O}^{n\times k}$ such that (6.7) holds, let $W\in\mathbb{O}^{k\times k}$ be an orthogonal polar factor of $\widehat{P}^{\operatorname{T}}\mathscr{H}(P)$, and set $\check{P}=\widehat{P}W_{1}$. Then $\check{P}^{\operatorname{T}}\mathscr{H}(P)=W_{1}^{\operatorname{T}}[\widehat{P}^{\operatorname{T}}\mathscr{H}(P)]\succeq 0$, and thus by Lemma B.9
@@
By the NPDo Ansatz, there exists $W_{2}\in\mathbb{O}^{k\times k}$ such that $\widetilde{P}=\check{P}W_{2}=\widehat{P}(W_{1}W_{2})\in\mathbb{P}$ and $f(\widetilde{P})\geq f(P)+(\omega/2)\eta$, verifying the NEPv Ansatz. ∎
@@
Suppose that the NEPv Ansatz holds, and let $P_{*}\in\mathbb{P}$ be a maximizer of (1.1). Then NEPv (6.3) holds for $P=P_{*}$ and the eigenvalues of $\Omega=\Omega_{*}:=P_{*}^{\operatorname{T}}H(P_{*})P_{*}$ consist of the first $k$ largest eigenvalues of $H(P_{*})$.
@@
$\max_{P\in\mathbb{O}^{n\times k}}\operatorname{tr}(P^{\operatorname{T}}H(P_{*})P).$ (6.13)
We claim $P_{*}$ is a maximizer of (6.13); otherwise there would be some $\widehat{P}\in\mathbb{O}^{n\times k}$ such that
@@ 6.2 SCF Iteration and Convergence
3:     solve SEP $H_{i}\widehat{P}^{(i)}=\widehat{P}^{(i)}\Omega_{i}$ for $\widehat{P}^{(i)}\in\mathbb{O}^{n\times k}$, an orthonormal basis matrix of the eigenspace associated with the first $k$ largest eigenvalues of $H_{i}$;
4:     calculate $Q_{i}\in\mathbb{O}^{k\times k}$ and let $P^{(i+1)}=\widehat{P}^{(i)}Q_{i}\in\mathbb{P}$, according to the NEPv Ansatz;
@@
At Line 4 it refers to the NEPv Ansatz for the calculation of $Q_{i}$. Exactly how it is computed depends on the structure of $f$ at hand. We commented on the similar issue for Algorithm 1 earlier. In the case of Example 6.1, $Q_{i}$ is taken to be an orthogonal polar factor of $(\widehat{P}^{(i)})^{\operatorname{T}}D$ to make $P^{(i+1)}\in\mathbb{O}_{D+}^{n\times k}$.
@@
Under the NEPv Ansatz, NEPv (6.3) is solvable, i.e., there exists $P\in\mathbb{O}^{n\times k}$ such that (6.3) holds and the eigenvalues of $\Omega=P^{\operatorname{T}}H(P)P$ are the $k$ largest ones of $H(P)$.
@@
    $\lambda_{k}(H(P_{*}Q))-\lambda_{k+1}(H(P_{*}Q))>0\quad\mbox{for any $Q\in\mathbb{O}^{k\times k}$},$ (6.16)
@@
    and if $f(P_{*})>f(P)$ for any $P\neq P_{*}$ and ${\cal R}(P)={\cal R}(P_{*})$, i.e., $f(P)$ has unique maximizer in the orbit $\{P_{*}Q\,:\,Q\in\mathbb{O}^{k\times k}\}$, then the entire sequence $\{P^{(i)}\}_{i=0}^{\infty}$ converges to $P_{*}$.
@@ 6.3 Acceleration by LOCG and Convergence
$W^{\operatorname{T}}\left(H(P)P-\frac{\partial f(P)}{\partial P}\right)=W^{\operatorname{T}}P\,\mathscr{M}(P)\quad\mbox{for $P=WZ$, $Z\in\mathbb{O}^{m\times k}$}.$
@@
$\big{(}W^{\operatorname{T}}H(WZ)W\big{)}Z-\frac{\partial\widetilde{f}(Z)}{\partial Z}=Z\,\widetilde{\mathscr{M}}(Z)\quad\mbox{for $Z\in\mathbb{O}^{m\times k}$},$ (6.18)
@@
$\widetilde{H}(Z)Z:=\big{(}W^{\operatorname{T}}H(WZ)W\big{)}Z=Z\widetilde{\Omega},\,\,Z\in\mathbb{O}^{m\times k}.$ (6.19)
By Theorem 6.1, we conclude that, in the case of (6.6), $Z\in\mathbb{O}^{m\times k}$ is a solution to the KKT condition (3.12) for the reduced problem (3.11b) if and only if it is a solution to NEPv (6.19) and ~$\widetilde{\mathscr{M}}(Z)$ is symmetric.
@@
Suppose that the NEPv Ansatz holds for $f(P)$ with $H(P)$, and let $\mathbb{Z}:=W^{\operatorname{T}}\mathbb{P}\subseteq\mathbb{O}^{m\times k}$. If $W\mathbb{Z}\subseteq\mathbb{P}$, then the NEPv Ansatz holds for $\widetilde{f}(Z)$ defined in (3.11b) with $\widetilde{H}(Z)=W^{\operatorname{T}}H(WZ)W$.
@@
Let $\widehat{Z}\in\mathbb{O}^{m\times k}$ and $Z\in\mathbb{Z}:=W^{\operatorname{T}}\mathbb{P}\subseteq\mathbb{O}^{m\times k}$ such that
@@
Set $P=WZ\in\mathbb{P}$ because of $W\mathbb{Z}\subseteq\mathbb{P}$ and $\widehat{P}=W\widehat{Z}\in\mathbb{O}^{n\times k}$. We get (6.7) from (6.20). By the NEPv Ansatz for $f$ with $H$, there exists $Q\in\mathbb{O}^{k\times k}$ such that $\widetilde{P}=\widehat{P}Q=W(\widehat{Z}Q)=:W\widetilde{Z}\in\mathbb{P}$ and $f(\widetilde{P})\geq f(P)+\omega\eta$, i.e.,
@@
3:     compute $W\in\mathbb{O}^{n\times m}$ such that ${\cal R}(W)={\cal R}([P^{(i)},\mathscr{R}(P^{(i)}),P^{(i-1)}])$ such that $P^{(i)}$ occupies the first $k$ columns of $W$;
@@ 7 Atomic Functions for NEPv
Armed with the general theoretical framework for the NEPv approach in section 6, in this section, we introduce the notion of atomic functions for NEPv, which serves as a singleton unit of function on $\mathbb{O}^{n\times k}_{\delta}$ for which the NEPv approach is guaranteed to work, and more importantly, the NEPv approach works for any convex composition of atomic functions, provided that some of the partial derivatives of the composing convex function are nonnegative.
In what follows, we first formulate two conditions that define the atomic functions and prove why the NEPv approach will work on atomic functions, and then we give concrete examples of atomic functions that encompass nearly all practical ones that are in use today, and we leave investigating how the NEPv approach will work on convex compositions of these atomic functions to section 8.
@@ 7.1 Conditions on Atomic Functions
Suppose that, for function $f$ defined on some neighborhood $\mathbb{O}^{n\times k}_{\delta}$ of $\mathbb{O}^{n\times k}$, we have already constructed an associated symmetric matrix-valued function $H(P)\in\mathbb{R}^{n\times n}$. We are interested in those that satisfy
$\operatorname{tr}(P^{\operatorname{T}}H(P)P)$ $=\underaccent{\bar}{\gamma}\,f(P)\quad\mbox{for $P\in\mathbb{P}\subseteq\mathbb{O}^{n\times k}$},$ (7.1a)
and given $P\in\mathbb{P}$ and $\widehat{P}\in\mathbb{O}^{n\times k}$, there exists $Q\in\mathbb{O}^{k\times k}$ such that $\widetilde{P}=\widehat{P}Q\in\mathbb{P}$ and
@@
where $\underaccent{\bar}{\alpha}>0,\,\underaccent{\bar}{\beta}\geq 0$, and $\underaccent{\bar}{\gamma}=\underaccent{\bar}{\alpha}+\underaccent{\bar}{\beta}$ are constants. Here a subset $\mathbb{P}\subseteq\mathbb{O}^{n\times k}$ is also involved.
@@
A function $f$ defined on some neighborhood $\mathbb{O}^{n\times k}_{\delta}$ of $\mathbb{O}^{n\times k}$ is an atomic function for NEPv if there are a symmetric matrix-valued function $H(P)\in\mathbb{R}^{n\times n}$ for $P\in\mathbb{O}^{n\times k}$ and constants $\underaccent{\bar}{\alpha}>0,\,\underaccent{\bar}{\beta}\geq 0$, and $\underaccent{\bar}{\gamma}=\underaccent{\bar}{\alpha}+\underaccent{\bar}{\beta}$ such that both conditions in (7.1) hold.
@@
$\operatorname{tr}(\widetilde{P}^{\operatorname{T}}H(P)\widetilde{P})\leq\underaccent{\bar}{\alpha}\,f(\widetilde{P})+\underaccent{\bar}{\beta}\,f(P)\quad\mbox{for $P,\,\widetilde{P}\in\mathbb{P}\subseteq\mathbb{O}^{n\times k}$},$ (7.1b)
without referring to an intermediate $\widehat{P}$. We claim that (7.1b) is stronger than (7.1b), assuming for any $\widehat{P}\in\mathbb{O}^{n\times k}$, there exists $Q\in\mathbb{O}^{k\times k}$ such that $\widetilde{P}=\widehat{P}Q\in\mathbb{P}$. Here is why. Suppose (7.1b) holds. Given any $\widehat{P}\in\mathbb{O}^{n\times k}$, let $\widetilde{P}=\widehat{P}Q\in\mathbb{P}$ for some $Q\in\mathbb{O}^{k\times k}$. Then by (7.1b) we have for any $P\in\mathbb{P}$
@@
yielding (7.1b). In view of this observation, in our later developments, we may verify (7.1b) if it can be verified.
@@
Assuming (4.1a), for $H(P)$ as given in (6.5), we have for $P\in\mathbb{P}\subseteq\mathbb{O}^{n\times k}$
@@
as was to be shown. Assume (4.1b). Let $\widehat{P}\in\mathbb{O}^{n\times k},\,P\in\mathbb{P}$ and let $W_{1}\in\mathbb{O}^{k\times k}$ be an orthogonal polar factor of $\widehat{P}^{\operatorname{T}}\mathscr{H}(P)$ and $\check{P}=\widehat{P}W_{1}$. Then we again have (6.12) and
@@
Now use (4.1b) to conclude that there is $W_{2}\in\mathbb{O}^{k\times k}$ such that $\widetilde{P}=\check{P}W_{2}=\widehat{P}(W_{1}W_{2})\in\mathbb{P}$ and
@@
Given function $f$ defined on $\mathbb{O}_{\delta}^{n\times k}$ and its associated symmetric $H(P)$ that satisfy (7.1), suppose $f(P)\geq 0$ for $P\in\mathbb{P}$. Let $g(P)=c[f(P)]^{s}$ where $c>0,\,s>1$, and let its associated symmetric matrix-valued function be $H_{g}(P)=cs\,[f(P)]^{s-1}H(P)$. Then
$\operatorname{tr}(P^{\operatorname{T}}H_{g}(P)P)$ $=s\underaccent{\bar}{\gamma}\,g(P)\quad\mbox{for $P\in\mathbb{P}\subseteq\mathbb{O}^{n\times k}$},$ (7.4a)
and given $P\in\mathbb{P}$ and $\widehat{P}\in\mathbb{O}^{n\times k}$, there exists $Q\in\mathbb{O}^{k\times k}$ such that $\widetilde{P}=\widehat{P}Q\in\mathbb{P}$ and
@@
Given $P\in\mathbb{P}\subseteq\mathbb{O}^{n\times k}$ and $\widehat{P}\in\mathbb{O}^{n\times k}$, suppose that (6.7) holds, i.e., $\operatorname{tr}(\widehat{P}^{\operatorname{T}}H(P)\widehat{P})\geq\operatorname{tr}(P^{\operatorname{T}}H(P)P)+\eta$. We have by (7.1)

### v4

@@ 6.1 The NEPv Ansatz
    It is tempting to stipulate $f(\widehat{P})\geq f(P)+\omega\eta$, but that is either false or just hard to prove for the one in Example 6.1 and some of those in Table 1 that involve $D$. Often in our algorithms to solve (1.1) iteratively, with $P$ being the current approximate maximizer, assuming the NEPv Ansatz, we naturally compute $\widehat{P}$ that maximizes $\operatorname{tr}(X^{\operatorname{T}}H(P)X)$ over $X\in{\rm St}(k,n)$. With that $\widehat{P}$, settling whether $f(\widehat{P})\geq f(P)+\omega\eta$ or not can be a hard task, as in Example 6.1 where the objective function involves $\operatorname{tr}((P^{\operatorname{T}}D)^{2})$.
@@
$H(P)$ conditions $\mathbb{P}$ by SEP $A$ none ${\rm St}(k,n)$ [20], [29, p.248] OLDA (6.4) with $\theta=1$, $D=0$ $B\succeq 0$, $s_{k}(B)>0$ ${\rm St}(k,n)$ (6.11) OCCA (6.4) with $\theta=1/2$, $A=0$ $B\succeq 0$, $s_{k}(B)>0$ ${\rm St}(k,n)_{D+}$ (6.11) $\Theta$TR (6.4) for case $10<\theta<$ 1 $B\succeq 0$, $s_{k}(B)>0$ ${\rm St}(k,n)_{D+}$ (6.11) $\operatorname{tr}(P^{\operatorname{T}}AP+P^{\operatorname{T}}D)\geq 0$ (6.4) for case $\theta=1$ $B\succeq 0$, $s_{k}(B)>0$ ${\rm St}(k,n)_{D+}$ (6.11) MBSub (6.4) with $\theta=0$ none ${\rm St}(k,n)_{D+}$ (6.11) SumCT (6.5) $A_{i}\succeq 0\,\forall i$ [66] Thm. 8.3 TrCP $2\sum_{i=1}^{N}\phi_{i}(\boldsymbol{x})A_{i}$ convex $\phi$ ${\rm St}(k,n)$ Expl. 8.1 UMDS $4\sum_{i=1}^{N}A_{i}PP^{\operatorname{T}}A_{i}$ $A_{i}\succeq 0\,\forall i$ ${\rm St}(k,n)$ Expl. 8.2 DFT $2A+2\sum_{i=1}^{n}\phi_{i}(\boldsymbol{x})\boldsymbol{e}_{i}\boldsymbol{e}_{i}^{\operatorname{T}}$ convex $\phi$ ${\rm St}(k,n)$ Expl. 8.1 * $\phi_{i}(\boldsymbol{x}):=\partial\phi(\boldsymbol{x})/\partial x_{i}$ for $\boldsymbol{x}=[x_{i}]$. $\Theta$TR for $\theta=0$ becomes MBSub.
As to the validity of the NEPv Ansatz on the objective functions in Table 1, it holds for all, except SumTR, under mild conditions on the constant matrices and function $\phi$. Table 4 provides the details on $H(P)$ and conditions under which the NEPv Ansatz holds, where the last column refers to places for justifications. We leave $\mathbb{P}$ unspecified for SumCT but refer it to [66] because it is more complicated to fit the space in the table. In fact, it is required that each $P_{i}$ falls in $\{X\in{\rm St}(k_{i},n)\,:\,X^{\operatorname{T}}D_{i}\succeq 0\}$. Later in subsection 8.2 we will argue that for SumCT it would be more efficient to go for the NPDo approach in Part I. For SumTR with $H(P)$ given by (6.2), $\mathbb{P}={\rm St}(k,n)$ and $Q=I_{k}$ and the NEPv Ansatz does not hold. This can be drawn from the counterexample, [76, Example 4.1], for which SCF diverges, but later we will show, under the NEPv Ansatz, SCF is guaranteed to converge!
@@
Let function $f$ be defined on some neighborhood ${\rm St}_{\delta}(k,n)$ of ${\rm St}(k,n)$ and let $H(P)$ be as in (6.5). Then the NPDo Ansatz implies the NEPv Ansatz.
@@
Suppose that the NPDo Ansatz holds. Given $\widehat{P}\in{\rm St}(k,n),\,P\in\mathbb{P}\subseteq{\rm St}(k,n)$ such that (6.7) holds, let $W\in{\rm St}(k,k)$ be an orthogonal polar factor of $\widehat{P}^{\operatorname{T}}\mathscr{H}(P)$, and set $\check{P}=\widehat{P}W_{1}$. Then $\check{P}^{\operatorname{T}}\mathscr{H}(P)=W_{1}^{\operatorname{T}}[\widehat{P}^{\operatorname{T}}\mathscr{H}(P)]\succeq 0$, and thus by Lemma B.9
@@
By the NPDo Ansatz, there exists $W_{2}\in{\rm St}(k,k)$ such that $\widetilde{P}=\check{P}W_{2}=\widehat{P}(W_{1}W_{2})\in\mathbb{P}$ and $f(\widetilde{P})\geq f(P)+(\omega/2)\eta$, verifying the NEPv Ansatz. ∎
@@
Let $P_{*}\in{\rm St}(k,n)$ be a maximizer of (1.1). Suppose that the NEPv Ansatz holds and $P_{*}\in\mathbb{P}$. Then NEPv (6.3) holds for $P=P_{*}$ and the eigenvalues of $\Omega=\Omega_{*}:=P_{*}^{\operatorname{T}}H(P_{*})P_{*}$ consist of the first $k$ largest eigenvalues of $H(P_{*})$.
@@
$\max_{P\in{\rm St}(k,n)}\operatorname{tr}(P^{\operatorname{T}}H(P_{*})P).$ (6.13)
We claim $P_{*}$ is a maximizer of (6.13); otherwise there would be some $\widehat{P}\in{\rm St}(k,n)$ such that
@@ 6.2 SCF Iteration and Convergence
3:     solve SEP $H_{i}\widehat{P}^{(i)}=\widehat{P}^{(i)}\Omega_{i}$ for $\widehat{P}^{(i)}\in{\rm St}(k,n)$, an orthonormal basis matrix of the eigenspace associated with the first $k$ largest eigenvalues of $H_{i}$;
4:     calculate $Q_{i}\in{\rm St}(k,k)$ and let $P^{(i+1)}=\widehat{P}^{(i)}Q_{i}\in\mathbb{P}$, according to the NEPv Ansatz;
@@
At Line 4 it refers to the NEPv Ansatz for the calculation of $Q_{i}$. Exactly how it is computed depends on the structure of $f$ at hand. We commented on the similar issue for Algorithm 1 earlier. In the case of Example 6.1, $Q_{i}$ is taken to be an orthogonal polar factor of $(\widehat{P}^{(i)})^{\operatorname{T}}D$ to make $P^{(i+1)}\in{\rm St}(k,n)_{D+}$.
@@
Under the NEPv Ansatz, NEPv (6.3) is solvable, i.e., there exists $P\in{\rm St}(k,n)$ such that (6.3) holds and the eigenvalues of $\Omega=P^{\operatorname{T}}H(P)P$ are the $k$ largest ones of $H(P)$.
@@
    $\lambda_{k}(H(P_{*}Q))-\lambda_{k+1}(H(P_{*}Q))>0\quad\mbox{for any $Q\in{\rm St}(k,k)$},$ (6.16)
@@
    and if $f(P_{*})>f(P)$ for any $P\neq P_{*}$ and ${\cal R}(P)={\cal R}(P_{*})$, i.e., $f(P)$ has unique maximizer in the orbit $\{P_{*}Q\,:\,Q\in{\rm St}(k,k)\}$, then the entire sequence $\{P^{(i)}\}_{i=0}^{\infty}$ converges to $P_{*}$.
@@ 6.3 Acceleration by LOCG and Convergence
$W^{\operatorname{T}}\left(H(P)P-\frac{\partial f(P)}{\partial P}\right)=W^{\operatorname{T}}P\,\mathscr{M}(P)\quad\mbox{for $P=WZ$, $Z\in{\rm St}(k,m)$}.$
@@
$\big{(}W^{\operatorname{T}}H(WZ)W\big{)}Z-\frac{\partial\widetilde{f}(Z)}{\partial Z}=Z\,\widetilde{\mathscr{M}}(Z)\quad\mbox{for $Z\in{\rm St}(k,m)$},$ (6.18)
@@
$\widetilde{H}(Z)Z:=\big{(}W^{\operatorname{T}}H(WZ)W\big{)}Z=Z\widetilde{\Omega},\,\,Z\in{\rm St}(k,m).$ (6.19)
By Theorem 6.1, we conclude that, in the case of (6.6), $Z\in{\rm St}(k,m)$ is a solution to the KKT condition (3.12) for the reduced problem (3.11b) if and only if it is a solution to NEPv (6.19) and ~$\widetilde{\mathscr{M}}(Z)$ is symmetric.
@@
Suppose that the NEPv Ansatz holds for $f(P)$ with $H(P)$, and let $\mathbb{Z}:=W^{\operatorname{T}}\mathbb{P}\subseteq{\rm St}(k,m)$. If $W\mathbb{Z}\subseteq\mathbb{P}$, then the NEPv Ansatz holds for $\widetilde{f}(Z)$ defined in (3.11b) with $\widetilde{H}(Z)=W^{\operatorname{T}}H(WZ)W$.
@@
Let $\widehat{Z}\in{\rm St}(k,m)$ and $Z\in\mathbb{Z}:=W^{\operatorname{T}}\mathbb{P}\subseteq{\rm St}(k,m)$ such that
@@
Set $P=WZ\in\mathbb{P}$ because of $W\mathbb{Z}\subseteq\mathbb{P}$ and $\widehat{P}=W\widehat{Z}\in{\rm St}(k,n)$. We get (6.7) from (6.20). By the NEPv Ansatz for $f$ with $H$, there exists $Q\in{\rm St}(k,k)$ such that $\widetilde{P}=\widehat{P}Q=W(\widehat{Z}Q)=:W\widetilde{Z}\in\mathbb{P}$ and $f(\widetilde{P})\geq f(P)+\omega\eta$, i.e.,
@@
3:     compute $W\in{\rm St}(m,n)$ such that ${\cal R}(W)={\cal R}([P^{(i)},\mathscr{R}(P^{(i)}),P^{(i-1)}])$ and $P^{(i)}$ occupies the first $k$ columns of $W$;
@@ 7 Atomic Functions for NEPv
Armed with the general theoretical framework for the NEPv approach in section 6, in this section, we introduce the notion of atomic functions for NEPv, which serves as a singleton unit of function on ${\rm St}_{\delta}(k,n)$ for which the NEPv approach is guaranteed to work for solving (1.1), and more importantly, the NEPv approach works on any convex composition of atomic functions, provided that some of the partial derivatives of the composing convex function are nonnegative.
In what follows, we first formulate two conditions that define atomic function and prove why the NEPv approach will work on the atomic functions, and then we give concrete examples of atomic functions that encompass nearly all practical ones that are in use today, and we leave investigating how the NEPv approach will work on convex compositions of these atomic functions to section 8.
@@ 7.1 Conditions on Atomic Functions
Suppose that, for function $f$ defined on some neighborhood ${\rm St}_{\delta}(k,n)$ of the Stiefel manifold ${\rm St}(k,n)$, we have already constructed an associated symmetric matrix-valued function $H(P)\in\mathbb{R}^{n\times n}$. We are interested in those that satisfy
$\operatorname{tr}(P^{\operatorname{T}}H(P)P)$ $=\underaccent{\bar}{\gamma}\,f(P)\quad\mbox{for $P\in\mathbb{P}\subseteq{\rm St}(k,n)$},$ (7.1a)
and given $P\in\mathbb{P}$ and $\widehat{P}\in{\rm St}(k,n)$, there exists $Q\in{\rm St}(k,k)$ such that $\widetilde{P}=\widehat{P}Q\in\mathbb{P}$ and
@@
where $\underaccent{\bar}{\alpha}>0,\,\underaccent{\bar}{\beta}\geq 0$, and $\underaccent{\bar}{\gamma}=\underaccent{\bar}{\alpha}+\underaccent{\bar}{\beta}$ are constants. Here a subset $\mathbb{P}\subseteq{\rm St}(k,n)$ is also involved.
@@
A function $f$ defined on some neighborhood ${\rm St}_{\delta}(k,n)$ of ${\rm St}(k,n)$ is an atomic function for NEPv if there are a symmetric matrix-valued function $H(P)\in\mathbb{R}^{n\times n}$ for $P\in{\rm St}(k,n)$ and constants $\underaccent{\bar}{\alpha}>0,\,\underaccent{\bar}{\beta}\geq 0$, and $\underaccent{\bar}{\gamma}=\underaccent{\bar}{\alpha}+\underaccent{\bar}{\beta}$ such that both conditions in (7.1) hold.
@@
$\operatorname{tr}(\widetilde{P}^{\operatorname{T}}H(P)\widetilde{P})\leq\underaccent{\bar}{\alpha}\,f(\widetilde{P})+\underaccent{\bar}{\beta}\,f(P)\quad\mbox{for $P,\,\widetilde{P}\in\mathbb{P}\subseteq{\rm St}(k,n)$},$ (7.1b)
without referring to an intermediate $\widehat{P}$. We claim that (7.1b) is stronger than (7.1b), assuming for any $\widehat{P}\in{\rm St}(k,n)$, there exists $Q\in{\rm St}(k,k)$ such that $\widetilde{P}=\widehat{P}Q\in\mathbb{P}$. Here is why. Suppose (7.1b) holds. Given any $\widehat{P}\in{\rm St}(k,n)$, let $\widetilde{P}=\widehat{P}Q\in\mathbb{P}$ for some $Q\in{\rm St}(k,k)$. Then by (7.1b) we have for any $P\in\mathbb{P}$
@@
yielding (7.1b). In view of this observation, in our later developments, we may verify (7.1b) directly if it can be verified.
@@
Assuming (4.1a), for $H(P)$ as given in (6.5), we have for $P\in\mathbb{P}\subseteq{\rm St}(k,n)$
@@
as was to be shown. Assume (4.1b). Let $\widehat{P}\in{\rm St}(k,n),\,P\in\mathbb{P}$ and let $W_{1}\in{\rm St}(k,k)$ be an orthogonal polar factor of $\widehat{P}^{\operatorname{T}}\mathscr{H}(P)$ and $\check{P}=\widehat{P}W_{1}$. Then we again have (6.12) and
@@
Now use (4.1b) to conclude that there is $W_{2}\in{\rm St}(k,k)$ such that $\widetilde{P}=\check{P}W_{2}=\widehat{P}(W_{1}W_{2})\in\mathbb{P}$ and
@@
Given function $f$ defined on ${\rm St}_{\delta}(k,n)$ and its associated symmetric $H(P)$ that satisfy (7.1), suppose $f(P)\geq 0$ for $P\in\mathbb{P}$. Let $g(P)=c[f(P)]^{s}$ where $c>0,\,s>1$, and let its associated symmetric matrix-valued function be $H_{g}(P)=cs\,[f(P)]^{s-1}H(P)$. Then
$\operatorname{tr}(P^{\operatorname{T}}H_{g}(P)P)$ $=s\underaccent{\bar}{\gamma}\,g(P)\quad\mbox{for $P\in\mathbb{P}\subseteq{\rm St}(k,n)$},$ (7.4a)
and given $P\in\mathbb{P}$ and $\widehat{P}\in{\rm St}(k,n)$, there exists $Q\in{\rm St}(k,k)$ such that $\widetilde{P}=\widehat{P}Q\in\mathbb{P}$ and
@@
Given $P\in\mathbb{P}\subseteq{\rm St}(k,n)$ and $\widehat{P}\in{\rm St}(k,n)$, suppose that (6.7) holds, i.e., $\operatorname{tr}(\widehat{P}^{\operatorname{T}}H(P)\widehat{P})\geq\operatorname{tr}(P^{\operatorname{T}}H(P)P)+\eta$. We have by (7.1)